OpenStudy (anonymous):

INCOME EXPENSES 1.)$37,025 -$34,492: Profit or Loss 2.)$35,175 -$32,490 :Profit or Loss 3.)$42,580 -$39,687 :Profit or Loss 4.)$50,032 -$45,510 :Profit or Loss 5.)$48,365 -$46,045 :Profit or Loss 6.)$42,115 -$38,518 :Profit or Loss 7.)$39,058 -$36,427:Profit or Loss 8.)$40,185 -$37,963:Profit or Loss 9.)$38,813 -$35,148:Profit or Loss 10.)$42,561 -$38,306:Profit or Loss 11.)$45,064 -$41,570:Profit or Loss

OpenStudy (anonymous):

@123baby

OpenStudy (anonymous):

ok so i just have to subtract?

OpenStudy (anonymous):

- means negative money

OpenStudy (anonymous):

@touseii45 @kelliegirl33 @phi @dmezzullo

OpenStudy (anonymous):

@CynthiaJune94 @cblrtopas

OpenStudy (anonymous):

What I need to know is if it's a loss or a profit

OpenStudy (anonymous):

seems like loss to me. they make 37,025 and the loss is 34,492. when you do a project you have to make at least twice as much as you spent. you use $600 to make a couch you need to make $1,000 to get you moneys worth to make more.

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

all of them?

OpenStudy (anonymous):

i would say ya, but you should get a second opinion

OpenStudy (anonymous):

id there more to the question?

OpenStudy (anonymous):

OKAY

OpenStudy (anonymous):

sorry caps lock

OpenStudy (anonymous):

it looks like OLAY

OpenStudy (anonymous):

i know right

OpenStudy (anonymous):

um...their all losses if you go by each individual one

OpenStudy (anonymous):

okay or(OLAY)0.o

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

and what do the equal @kelliegirl33

OpenStudy (anonymous):

you MINUS the problems

OpenStudy (anonymous):

Are they each separate problems or is it just one large problem ?

OpenStudy (anonymous):

so they are each separate problems...correct ? Or do I add ALL of the income and ALL of the expenses and total them up at the end ?

OpenStudy (anonymous):

separate!

OpenStudy (anonymous):

1. 37025 - 34492 = 2533 loss 2. 35175 - 32490 = 2685 loss 3. 42580 - 39687 = 2893 loss 4. 50032 - 45510 = 4522 loss 5. 48365 - 46045 = 2320 loss 6. 42115 - 38518 = 3597 loss 7. 39058 - 36427 = 2631 loss 8. 40185 - 37963 = 2222 loss 9. 38813 - 35148 = 3665 loss 10. 42561 - 38306 = 4255 loss 11. 45064 - 41570 = 3494 loss

OpenStudy (anonymous):

okay thank you now can you a line graph now @kelliegirl33

OpenStudy (anonymous):

I am not good at graphs

OpenStudy (anonymous):

to me, they look like the income and expenses for one year and you are asked to find if there was an annual profit or loss. everytime, we notice that income > expenses. so, think about it. If you get an allowance from parents for $10 every month and you spend $7, then you SAVE $3 every month. So, it is a profitable venture. In your question, you just provide a table but no actual question. So, there is no way to help you. You should provide the complete problem statement if you want a proper help.

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

they are all profit

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

bc your income is what you make expense is what you spend

OpenStudy (anonymous):

say you own a business you spend 50,000 to open after the first month your income is 60,000 you make a 10,000 profit

OpenStudy (anonymous):

Yes I know that but my friend said they were all Loss? *confused*

OpenStudy (anonymous):

Maybe can you show me with one of my questions. *they are all separate*

OpenStudy (anonymous):

Not one big question.

OpenStudy (anonymous):

okay income=$37025-$34,492=$2533

OpenStudy (anonymous):

On all of the EXPENSES the - is a negative sign okay.

OpenStudy (anonymous):

that just shows how much money you spent but you still made more than you spent

OpenStudy (anonymous):